# When Ads Become Profiles: Uncovering the Invisible Risk of Web Advertising at Scale with LLMs

### v2

@@ 4.3. User-Level User Profile Reconstruction
While session-level analysis captures short-term context, a user’s full ad exposure history offers a richer and more stable signal. To leverage this, we construct a user-level profile by aggregating the reasoning summaries derived in the previous stage. Let $R_{u}=(r_{s_{1}},r_{s_{2}},\dots,r_{s_{m}})$ denote the chronological sequence of reasoning summaries generated for user $u$ across $m$ sessions. These summaries are concatenated to form a longitudinal narrative, representing a condensed history of the inferred demographic signals over the entire two-year observation period. The final profiling is performed by the model $\mathcal{M}$, conditioned on a synthesis prompt $\mathcal{P}_{\text{profile user}}$, to predict the final attribute set $\hat{\mathcal{A}}_{u}$:
(4) $\hat{\mathcal{A}}_{u}=\mathcal{M}(R_{u}\mid\mathcal{P}_{\text{profile user}})$
This hierarchical aggregation allows the adversary to synthesise cumulative evidence (i.e., aggregate semantic signals) and detect temporal patterns (e.g., shifts in interests or trends) from the full reasoning summaries of an user $R_{u}$. Crucially, this approach remains computationally tractable by processing condensed summaries rather than the prohibitive raw ad stream $\mathcal{S}_{u}$.
@@ 5.1. Dataset
To empirically evaluate our threat model, we utilise a large-scale, longitudinal dataset sourced from a data donation project. The project recruits volunteer participants from the Australian public to donate data about the advertisements they encounter on Facebook via a custom-built, privacy-preserving browser plugin. Crucially, this collection mechanism mirrors the exact vantage point of the malicious browser extension described in our threat model (Section 3) it passively captures the stream of sponsored posts directly from the user’s feed as they browse the website. Participants joined by installing the plugin and completing a detailed demographic questionnaire, which provides the high-fidelity ground truth necessary to benchmark adversarial inference accuracy. The collected data for each ad includes the ad creative (image and text), associated metadata, and a link to the de-identified user profile. This methodology provides a unique, user-centric view of the ad ecosystem, capturing the personalised ads that are otherwise inaccessible to public scrutiny via scraping or ad libraries. The full dataset made available to us comprises over 700,000 ad observations collected between 2021 and 2023 including demographic information from over 2,000 Australian Facebook users. This extensive, real-world dataset allows us to conduct a robust, large-scale study of the potential for user profiling attack by LLMs under realistic conditions.
@@ 5.2. Preprocessing
We compute a session threshold $\theta$ by applying a kernel density estimator (KDE) over the log-transformed inter-ad time intervals. This process yields a global threshold of $\theta=389$ seconds (refer to Appendix E for the detailed preprocessing details). Thus, two consecutive ads $t_{i},t_{i+1}$ are considered to belong to different sessions if $t_{i+1}-t_{i}>\theta$.
This data-driven approach allows us to transform the continuous variable of time gaps into a principled categorical distinction (same session vs. new session) based on real user behavioural patterns observed in the data. Following temporal segmentation, we apply additional filtering steps to ensure data quality and modelling tractability: (1) Source filtering: We retain only ad impressions sourced from Facebook, excluding other platforms and modalities such as video-only content. (2) Session length bounds: We discard sessions with fewer than 3 or more than 50 ad impressions to eliminate sparse sessions and abnormally long sequences. This range aligns with recent studies on LLM-based user profiling and prediction on sequential data (feng2025agentmove; wongso2024genup). (3) User filtering: We exclude users who contributed fewer than 3 ad session, as such short histories are insufficient for meaningful profiling.
After filtering, the dataset contains: $891$ unique users, 63,86463{,}864 total ad sessions and 435,314435{,}314 total ad impressions. To further prepare the ad text for subsequent processing, we remove all HTML elements and markup to produce clean plain-text representations. The demographic composition of the final filtered user cohort is detailed in Table 4. This table provides the operational definitions for all demographic categories, such as age ranges (older or younger), income brackets and education levels (higher or lower), used throughout our subsequent analyses. All monetary values, including income brackets, are reported in Australian Dollars (AUD). The distribution of ad content across IAB categories, which forms a core feature for our subsequent analyses, is visualised in Figure 4. These distributions provide the foundational context for our user profile reconstruction experiments.
@@ 5.3. User Profile Reconstruction
Experimental Setup. We evaluate model performance using Accuracy and Macro F1-score, ensuring robust assessment even in the presence of class imbalance. For ordinal attributes such as Age and Income, where exact classification is overly punitive and “near misses” retain profiling utility, we additionally report Mean Absolute Error (MAE). This measures the average index distance between the predicted and ground-truth bins. To allow for comparison across attributes with differing granularities (e.g., 7 age brackets vs. 12 income brackets), we also report Normalised MAE (NMAE). Complete model hyperparameters and configurations are detailed in Appendix B.
@@ 5.3.1. Session-Level Reconstruction
We first conduct our analysis at the most granular level: the user session. A user session is defined as the sequence of ads a user is exposed to within a single, continuous period of platform usage. This session-level analysis allows us to test the LLM’s inference capabilities on short, contextually coherent ad sequences. To ensure data quality and meaningful classification, we excluded responses where users selected “prefer not to say” for any given attribute. We also excluded the “other” category for gender due to its small sample size (N=6 users), which is insufficient for a reliable model evaluation. We then tasked Gemini 2.0 Flash, a state-of-the-art multimodal LLM, with performing zero-shot classification for six demographic attributes across the remaining 63,864 user sessions from our dataset.
@@
Human vs. AI Evaluation. To contextualise these LLMs capabilities, we conducted a human evaluation on randomly sampled 100 user sessions, stratified to match the demographic distribution of our dataset as detailed in Table 4. We recruited six human annotators (2 female, 4 male) with diverse professional backgrounds, including legal and privacy expertise, computer science, and human resources. Each annotator was tasked with inferring the six demographic attributes for each of the 100 sessions, based solely on the same ad content shown to the LLM. To best capture intuitive human inference in a realistic and tractable manner, annotators were shown the raw, human-readable visual images of the ads within each session, as processing the full structured textual captions, categories, key entities from all ads in 100 sessions would be cognitively overwhelming. In contrast, the LLM received the structured textual input generated by our multimodal understanding pipeline (as described in Section 4.1) as in our session-level analysis. This design allows for an valid comparison between intuitive human inference and systematic algorithmic inference. The results are detailed in Table 5.3 and visualised in Figure 3.
@@ 5.3.2. User-Level Reconstruction
Building on the hierarchical aggregation pipeline defined in Section 4.3, we reconstruct longitudinal user profiles by synthesizing reasoning summaries across all sessions. This holistic approach allows the model to leverage cumulative evidence and temporal shifts in ad exposure over the user’s entire history. For this more rigorous user-level evaluation, we introduce two stronger baseline models derived from Australian census data (ABS2021Census), which reflect the prior demographic distribution of the population. The Prior-Mode baseline always predicts the majority class for each attribute. The Prior-Sampling baseline provides a more robust chance-level comparison by making predictions through random sampling from the census distribution; we report the mean and standard deviation of its performance over 1,000 independent runs in Table 3. To ensure a fair comparison, we harmonised our dataset with the census data. For all attributes, we excluded responses where users selected “prefer not to say.” We excluded gender-diverse respondents from this analysis because the census prior reports only male and female categories. For education, we aligned with the higher education census distribution by merging all non-degree categories in our data into a single “No Degree” class. Census data for individual income, reported weekly, was annualised and aligned perfectly with our income brackets. For employment, we merged our two “unemployed” categories to match the census, which does not distinguish between looking for work or not; we also excluded “retired” individuals from this specific attribute’s evaluation, as the census employment data pertains only to the labour force. Detailed demographic statistics are based on the Australian Census (ABS2021Census) and the Australian Federal Election Study (cameron20222022). We evaluate three versions of LLM: the standard Gemini 2.0 flash model, an augmented version, Gemini, which was provided with a prompt about the Australian cultural context and a a “Random Control” group (where ad session order within an user is shuffled). This augmentation is designed to create a more equitable comparison with our census-derived baselines. Since the Prior-Sampling baseline inherently incorporates “Australian context” by drawing from national demographic data, providing the LLM with similar high-level contextual information ensures that our evaluation more accurately assesses the model’s ability to infer information from ad content, rather than simply penalising it for a lack of geographical and cultural knowledge.
@@
To isolate the value of sequential information, we compare the standard model against a “Random Control” group where the chronological order of ads within sessions was shuffled. The results indicate that temporal order contributes differently across attributes. For Age and Gender, preserving the chronological sequence yields noticeable gains (e.g., Age accuracy improves from 38.00% to 41.18%, and MAE decreases to 0.85), suggesting that temporal evolution in ad delivery captures signals related to life stages or identity stability. However, for socio-economic attributes like Income, Education, and Employment, performance remains comparable between the sequential and shuffled inputs. This implies that for these traits, the cumulative semantic content: the specific collection of ads a user sees, serves as the primary predictive signal, regardless of the specific viewing order.
@@ 6. Discussion
The Shift to Passive Privacy Risk. Our findings mark a critical shift in the understanding of digital privacy. The high accuracy of our LLM-based reconstruction, which consistently outperforms census priors and matches human intuition, confirms that generative AI can now systematically decode granular demographic signals inherently encoded in ad-delivery algorithms. Furthermore, our analysis of ordinal attributes (Age, Income) reveals that even when the model’s exact predictions fail, they remain “directionally correct” (low MAE), placing users in semantically accurate categories (e.g., “young”, “low-income”). The democratisation of multimodal LLMs transforms this from a theoretical risk into an immediate operational threat. Unlike traditional profiling that required collecting massive labeled datasets to train bespoke classifiers (bi2013inferring; kosinski2013private; schwartz2013personality; hinds2018demographic; tricomi2024spotify), our zero-shot approach demonstrates that adversaries can now leverage off-the-shelf models to perform scalable profiling with minimal technical expertise. While prior research demonstrated that LLMs can infer attributes from user content such as Reddit posts (staab24beyond; tomekcce2024private) or personal photo albums (liu2025eye), our work reveals that passive consumption of algorithmic content constitutes an equally potent leakage vector. This capability is particularly dangerous within the browser extension ecosystem (singh2025study), where benign-looking tools (e.g., ad blockers) possess legitimate permissions to read page content. An adversary can thus deploy a “silent” profiler that bypasses platform targeting restrictions (e.g., on political affiliation) by reverse-engineering the platform’s own optimisation logic locally on the user’s device. This vector exploits a critical gap in user understanding: users typically focus privacy concerns on invisible trackers and cookies, rarely perceiving the visible ad creative itself as a leakage vector (yao2017folk). This creates a privacy dilemma: users can choose not to post content, but they cannot easily opt out of the ad ecosystem, and they are often unaware that even passive exposure can be harvested to reconstruct their identity. Taken together, these findings suggest that current privacy policy may underestimate risks by focusing on exact identification rather than semantic profiling.
@@
Limitations. Our study has limitations that contextualise these findings. First, our dataset is drawn from a self-selected cohort of Australian Facebook users via a desktop browser plugin. This sample may not fully generalise to mobile-only populations or other cultural and regulatory contexts (kaushik2024cross). Second, our analysis is limited to Facebook; other ecosystems like TikTok or open programmatic web advertising involve different ad formats and delivery logics that may yield different leakage patterns. Finally, our reconstruction pipeline separates visual captioning from reasoning to manage token costs. While effective, a end-to-end multimodal model might capture even subtler visual cues, potentially yielding higher inference accuracy and greater privacy risk than reported here. Future work should extend this framework to cross-platform environments and explore adversarial defenses against LLM-based profiling.
@@ 7. Conclusion
This work reveals a critical blind spot in Web privacy: the latent leakage of user private attributes through passive exposure to algorithmic advertising. By leveraging multimodal LLMs as adversarial inference engines, we demonstrate ad streams serve as high-fidelity digital footprints, enabling accurate reconstruction of private attributes, often surpassing human social perception. Our findings confirm that this threat is operational and scalable: actionable profiles can be derived from short observation windows without long-term tracking, and even “near-miss” predictions for age and income yield semantically invasive insights. As off-the-shelf AI tools democratise this capability, they increase the risk that benign browser extensions could be repurposed to circumvent platform targeting restrictions. Addressing this risk requires moving beyond current regulatory frameworks to recognise and govern the latent semantic signals embedded in the content users passively consume.
@@ Acknowledgements.
This research includes computations using the Wolfpack computational cluster, supported by the School of Computer Science and Engineering at UNSW Sydney. We also acknowledge support from the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005). We would like to thank Abdul Obeid for suggesting the idea of using a KDE-based approach for defining session thresholds. We further thank Baihe Zhang, Abdul Obeid, Zechen Li, Wilson Wongso, Yixuan Liu for their hard work on annotation, which greatly contributed to the human evaluation.

### v3

@@ 4.3. User-Level User Profile Reconstruction
While session-level analysis captures short-term context, a user’s full ad exposure history offers a richer and more stable signal. To leverage this, we construct a user-level profile by aggregating the reasoning summaries derived in the previous stage. Let $R_{u}=(r_{s_{1}},r_{s_{2}},\dots,r_{s_{m}})$ denote the chronological sequence of reasoning summaries generated for user $u$ across $m$ sessions. These summaries are concatenated to form a longitudinal narrative, representing a condensed history of the inferred demographic signals over the entire two-year observation period. The final profiling is performed by the model $\mathcal{M}$, conditioned on a synthesis prompt $\mathcal{P}_{\text{profile user}}$, to predict the final attribute set $\hat{\mathcal{A}}_{u}$: $\hat{\mathcal{A}}_{u}=\mathcal{M}(R_{u}\mid\mathcal{P}_{\text{profile user}})$. This hierarchical aggregation allows the adversary to synthesise cumulative evidence (i.e., aggregate semantic signals) and detect temporal patterns (e.g., shifts in interests or trends) from the full reasoning summaries of a user $R_{u}$. Crucially, this approach remains computationally tractable by processing condensed summaries rather than the prohibitive raw ad stream $\mathcal{S}_{u}$.
@@ 5.1. Dataset
To empirically evaluate our threat model, we utilise a large-scale, longitudinal dataset sourced from the Australian Ad Observatory, a citizen science project run by the ARC Centre of Excellence for Automated Decision-Making and Society (ADM+S) (Angus et al., 2024a, b). The project recruits volunteer participants from the Australian public to donate data about the advertisements they encounter on Facebook via a custom-built, privacy-preserving browser plugin (Angus et al., 2024a, b). Crucially, this collection mechanism mirrors the exact vantage point of the malicious browser extension described in our threat model (Section 3), it passively captures the stream of sponsored posts directly from the user’s feed as they browse the website. Participants joined by installing the plugin and completing a detailed demographic questionnaire, which provides the high-fidelity ground truth necessary to benchmark adversarial inference accuracy. The collected data for each ad includes the ad creative (image and text), associated metadata, and a link to the de-identified user profile. This methodology provides a unique, user-centric view of the ad ecosystem, capturing the personalised ads that are otherwise inaccessible to public scrutiny via scraping or ad libraries. The full dataset made available to us comprises over 700,000 ad observations collected between 2021 and 2023 including demographic information from over 2,000 Australian Facebook users. This extensive, real-world dataset allows us to conduct a robust, large-scale study of the potential for user profiling attack by LLMs under realistic conditions.
@@ 5.2. Preprocessing
We compute a session threshold $\theta$ by applying a kernel density estimator (KDE) over the log-transformed inter-ad time intervals. This process yields a global threshold of $\theta=389$ seconds (refer to Appendix E for the detailed preprocessing details). Thus, two consecutive ads $t_{i},t_{i+1}$ are considered to belong to different sessions if $t_{i+1}-t_{i}>\theta$. This data-driven approach allows us to transform the continuous variable of time gaps into a principled categorical distinction (same session vs. new session) based on real user behavioural patterns observed in the data. Following temporal segmentation, we apply additional filtering steps to ensure data quality and modelling tractability: (1) Source filtering: We retain only ad impressions sourced from Facebook, excluding other platforms and modalities such as video-only content. (2) Session length bounds: We discard sessions with fewer than 3 or more than 50 ad impressions to eliminate sparse sessions and abnormally long sequences. This range aligns with recent studies on LLM-based user profiling and prediction on sequential data (Feng et al., 2025; Wongso et al., 2025). (3) User filtering: We exclude users who contributed fewer than 3 ad sessions, as such short histories are insufficient for meaningful profiling.
After filtering, the dataset contains: $891$ unique users, 63,86463{,}864 total ad sessions and 435,314435{,}314 total ad impressions. To prepare the ad text for subsequent processing, we remove all HTML elements and markup to produce clean plain-text. The demographic composition of the final filtered user cohort is detailed in Table 4. This table provides the operational definitions for all demographic categories, such as age ranges (older or younger), income brackets and education levels (higher or lower), used throughout our analyses. All monetary values, including income brackets, are reported in Australian Dollars (AUD). The distribution of ad content across IAB categories, which forms a core feature for our subsequent analyses, is visualised in Figure 4. These distributions provide the foundational context for our user profile reconstruction experiments.
@@ 5.3. User Profile Reconstruction
Experimental Setup. We evaluate performance using Accuracy and Macro F1-score, ensuring robustness under class imbalance. For ordinal attributes such as Age and Income, where exact classification is overly punitive and “near misses” retain profiling utility, we additionally report Mean Absolute Error (MAE). This measures the average index distance between the predicted and ground-truth bins. To allow for comparison across attributes with differing granularities (e.g., 7 age brackets vs. 12 income brackets), we also report Normalised MAE (NMAE). Complete model hyperparameters and configurations are detailed in Appendix B.
@@ 5.3.1. Session-Level Reconstruction
We first conduct our analysis at the most granular level: the user session, defined as the sequence of ads a user is exposed to within a single, continuous period of platform usage. This session-level analysis tests the LLM’s inference capabilities on short, contextually coherent ad sequences. To ensure data quality and meaningful classification, we excluded responses where users selected “prefer not to say” for any given attribute. We also excluded the “other” category for gender due to its small sample size (N=6), which is insufficient for a reliable model evaluation. We then tasked Gemini 2.0 Flash, a state-of-the-art multimodal LLM, with zero-shot classification for six demographic attributes across the remaining 63,864 user sessions from our dataset.
@@
Human vs. AI Evaluation. To contextualise these LLMs’ capabilities, we conducted a human evaluation on randomly sampled 100 user sessions, stratified to match the demographic distribution of our dataset as detailed in Table 4. We recruited six human annotators (2 female, 4 male) with diverse professional backgrounds, including legal and privacy expertise, computer science, and human resources. Each annotator was tasked with inferring the six demographic attributes for each of the 100 sessions, based solely on the same ad content shown to the LLM. To best capture intuitive human inference in a realistic and tractable manner, annotators were shown the raw, human-readable visual images of the ads within each session, as processing the full structured textual captions, categories, key entities from all ads in 100 sessions would be cognitively overwhelming. In contrast, the LLM received the structured textual input generated by our multimodal understanding pipeline (as described in Section 4.1) as in our session-level analysis. This design allows for a valid comparison between intuitive human inference and systematic algorithmic inference. The results are detailed in Table 2 and visualised in Figure 3.
@@ 5.3.2. User-Level Reconstruction
Building on the hierarchical aggregation pipeline defined in Section 4.3, we reconstruct longitudinal user profiles by synthesizing reasoning summaries across all sessions. This holistic approach allows the model to leverage cumulative evidence and temporal shifts in ad exposure over the user’s entire history. For this more rigorous user-level evaluation, we introduce two stronger baseline models derived from Australian census data (Australian Bureau of Statistics, 2021), which reflect the prior demographic distribution of the population. The Prior-Mode baseline always predicts the majority class for each attribute. The Prior-Sampling baseline provides a more robust chance-level comparison by making predictions through random sampling from the census distribution; we report the mean and standard deviation of its performance over 1,000 independent runs in Table 3. To ensure a fair comparison, we harmonised our dataset with the census data. For all attributes, we excluded responses where users selected “prefer not to say.” We excluded gender-diverse respondents from this analysis because the census prior reports only male and female categories. For education, we aligned with the higher education census distribution by merging all non-degree categories in our data into a single “No Degree” class. Census data for individual income, reported weekly, was annualised and aligned perfectly with our income brackets. For employment, we merged our two “unemployed” categories to match the census, which does not distinguish between looking for work or not; we also excluded “retired” individuals from this specific attribute’s evaluation, as the census employment data pertains only to the labour force. Detailed demographic statistics are based on the Australian Census (Australian Bureau of Statistics, 2021) and the Australian Federal Election Study (Cameron et al., 2022). We evaluate three versions of LLM: the standard Gemini 2.0 flash model, an augmented version, Gemini, which was provided with a prompt about the Australian cultural context and a “Random Control” group (where ad session order within a user is shuffled). This augmentation is designed to create a more equitable comparison with our census-derived baselines. Since the Prior-Sampling baseline inherently incorporates “Australian context” by drawing from national demographic data, providing the LLM with similar high-level contextual information ensures that our evaluation more accurately assesses the model’s ability to infer information from ad content, rather than simply penalising it for a lack of geographical and cultural knowledge.
@@
To isolate sequential information, we compare the standard model against a “Random Control” group where the order of ads within sessions was shuffled. The results indicate that temporal order contributes differently across attributes. For Age and Gender, preserving the chronological sequence yields gains (e.g., Age accuracy improves from 38.00% to 41.18%, and MAE decreases to 0.85), suggesting that temporal evolution in ad delivery captures signals related to life stages or identity stability. However, for socio-economic attributes like Income, Education, and Employment, performance remains comparable between the sequential and shuffled inputs. This implies that for these traits, the cumulative semantic content: the specific collection of ads a user sees, serves as the primary predictive signal, regardless of the specific viewing order.
@@ 6. Discussion
The Shift to Passive Privacy Risk. Our findings mark a shift in digital privacy understanding. Our LLM-based reconstruction, which outperforms census priors and matches human intuition, confirms that generative AI can systematically decode granular demographic signals encoded in ad-delivery algorithms. Furthermore, our analysis of ordinal attributes (Age, Income) reveals that when exact predictions fail, they remain “directionally correct” (low MAE), placing users in semantically accurate categories (e.g., “young”, “low-income”). The democratisation of multimodal LLMs transforms this from a theoretical risk into an immediate operational threat. Unlike traditional profiling that required collecting massive labeled datasets to train bespoke classifiers (Bi et al., 2013; Kosinski et al., 2013; Schwartz et al., 2013; Hinds and Joinson, 2018; Tricomi et al., 2024), our zero-shot approach demonstrates that adversaries can now leverage off-the-shelf models to perform scalable profiling with minimal technical expertise. While prior research demonstrated that LLMs can infer attributes from user content such as Reddit posts (Staab et al., 2024; Tömekçe et al., 2024) or personal photo albums (Liu et al., 2025), our work reveals that passive consumption of algorithmic content constitutes a comparable leakage vector. This capability is particularly dangerous within the browser extension ecosystem (Singh et al., 2025), where benign-looking tools (e.g., ad blockers) possess legitimate permissions to read page content. An adversary can thus deploy a “silent” profiler that bypasses platform targeting restrictions (e.g., on political affiliation) by reverse-engineering the platform’s own optimisation logic locally on the user’s device. This vector exploits a critical gap in user understanding: users typically focus privacy concerns on invisible trackers and cookies, rarely perceiving the visible ad creative itself as a leakage vector (Yao et al., 2017). This creates a privacy dilemma: users can choose not to post content, but they cannot easily opt out of the ad ecosystem, and they are often unaware that even passive exposure can be harvested to reconstruct their identity. Taken together, these findings suggest that current privacy policy may underestimate risks by focusing on exact identification rather than semantic profiling.
@@
Limitations. We note several limitations. First, our dataset is drawn from a self-selected cohort of Australian Facebook users via a desktop browser plugin. This sample may not fully generalize to mobile-only populations or other cultural and regulatory contexts (Kaushik et al., 2024). Second, our analysis is limited to Facebook; other ecosystems like TikTok or open programmatic web advertising involve different ad formats and delivery logics that may yield different leakage patterns. Finally, our reconstruction pipeline separates visual captioning from reasoning to manage token costs. While effective, a fully end-to-end multimodal model might capture even subtler visual cues, potentially yielding higher inference accuracy and greater privacy risk than reported here. Future work should extend this framework to cross-platform environments and explore adversarial defenses against LLM-based profiling.
@@ 7. Conclusion
This work reveals a critical blind spot in Web privacy: the latent leakage of user private attributes through passive exposure to algorithmic advertising. By leveraging multimodal LLMs as adversarial inference engines, we demonstrate that ad streams serve as high-fidelity digital footprints, enabling the accurate reconstruction of private attributes, often surpassing human social perception. Our findings confirm that this threat is operational and scalable: actionable profiles can be derived from short observation windows without long-term tracking, and “near-miss” predictions for age and income yield semantically invasive insights. As off-the-shelf AI tools democratise this capability, they increase the risk that benign browser extensions could be repurposed to circumvent platform targeting restrictions. Addressing this risk requires moving beyond current regulatory frameworks to recognise and govern the latent semantic signals embedded in the content users passively consume.
@@ Acknowledgements.
This research includes computations using the Wolfpack computational cluster, supported by the School of Computer Science and Engineering at UNSW Sydney. We acknowledge support from the Australian Research Council (ARC) Centre of Excellence for Automated Decision-Making and Society (ADM+S; CE200100005), which hosts the Australian Ad Observatory project, and thank the Ad Observatory team for their work and support. We thank Abdul Obeid for suggesting the idea of using a KDE-based approach for defining session thresholds. We further thank our annotators for their hard work, which greatly contributed to the human evaluation.